OpenStudy (anonymous):

The atmospheric pressure P in pounds per square inch (psi) is given by P = 14.7 e−0.21 a where a is the altitude above sea level (in miles). If a city has an atmospheric pressure of 13.17 psi, what is its altitude? (Recall that 1 mi = 5,280 ft. Round your answer to the nearest foot.) ft

ganeshie8 (ganeshie8):

\(\large P = 14.7 e^{−0.21} a \)

ganeshie8 (ganeshie8):

you're given \(P = 13.17\), plug it in and solve \(a\)

ganeshie8 (ganeshie8):

\(\large 13.17 = 14.7 e^{−0.21} a \)

ganeshie8 (ganeshie8):

can u solve \(a\) ?

OpenStudy (anonymous):

I'm sorry I was writing a a paper

OpenStudy (anonymous):

he atmospheric pressure P in pounds per square inch (psi) is given by P = 14.7 (e−0.21 a) where a is the altitude above sea level (in miles). If a city has an atmospheric pressure of 11.21 psi, what is its altitude? (Recall that 1 mi = 5,280 ft. Round your answer to the nearest foot.) -------------- 11.21 = 14.7*e^-0.21a ------------ e^(-0.21a) = 0.7626 ------- Take the natural log of both sides to get: -0.21a = -0.2710 ---- a = 1.2906 miles --- a = 1.2906*5280 = 6814 feet

OpenStudy (anonymous):

i found a question similar to that

OpenStudy (anonymous):

try using it to solve your question!

OpenStudy (anonymous):

OK THX